OpenStudy (arivillafana):

I'll FAN and MEDAL What is the Slope, Distance, and Midpoint of the following: A(0,5) B(7,-5) A(4,-5) B(0,-5)

OpenStudy (qwertty123):

Hey I will show you the Midpoint first.

OpenStudy (arivillafana):

awesome !

OpenStudy (anonymous):

Slope= -10/7 Midpoint=3 1/2 Distant:-6 Slope= 0/4 Midpoint=1,-5 Distant: -36

OpenStudy (qwertty123):

The midpoint between points A and B is : M = ( 72 , 0 ) Explanation To find midpoint between points A(xA,yA) and B(xB,yB), we use formula: M=(xA+xB2,yA+yB2) In this example: xA=0 , yA=5 , xB=7 and yB=−5 so: MM=(0+72,5+(−5)2)=( 72 , 0 )

OpenStudy (arivillafana):

which order is that in ?

OpenStudy (qwertty123):

Wow @Keyairra22 You need to explain also not just give direct answers!

OpenStudy (arivillafana):

would you help me with the bottom two ? @Qwertty123

OpenStudy (anonymous):

@Qwertty123 don't worry about me